I was wondering if anyone could help me with a mixture
A chemist needs to prepare 12oz of a 9% hydrochloric acid solution. Find the amount of a 4% solution and the amount of a 12% solution she should mix to get this solution.
Thanks for the help

Found 2 solutions by solver91311, vleith:
Answer by solver91311(12126)


The total amount of solution has to be 12 ounces, so let the amount of 4% solution be x ounces and the amount of 12% solution be 12 - x ounces.

4% of the x ounces of 4% solution is pure acid. 12% of the 12 - x ounces of 12% solution is pure acid. When you add these two together, you get something that is 9% times 12 ounces pure acid. Symbolically,



Solve for x to get the amount of 4% solution. Subtract this result from 12 to get the amount of 12% solution. Hint: This will be much easier to avoid errors if you first multiply both sides of the equation by 100 to get rid of the decimal fractions.

John


Answer by vleith(2517)

The total amount of acid is 12*0.09 = 1.08
Now you need to decide how much 4% needs to be mixed with 12% to yield that same amount of acid.
Let x be the amount of 4%. Let y be the amount of 12%.
You know that x + y = 12
So y = 12-x
TotalAcid = 0.04x + 0.12y
1.08 = 0.04x + 0.12(12-x)
1.08 = 0.04x + 1.44 - 0.12x
-0.36 = - 0.08x
4.5 = x
That means y = 12-4.5 = 7.5
So we need 7.5oz og 12% and 4.5oz of 4%
Make sure to check your answer
Does 4.5*0.04 + 7.5*0.12 = 12*0.09???